Two thin(small angled)prism are combined to produce dispersion without deviation. One prism has angle 5A^° and refractive index 1.56.if the other prism has refractive index 1.7,what is its angle?
1
Expert's answer
2019-10-22T10:17:06-0400

Angle of Prism


We need to find the Angle of the second prism


Solution:


We know the formula to find the deviation produced by a Prism, that is



"d = (\mu - 1) A"

In this,


d = Deviation produced by the Prism


"\mu =" Refractive index of the Prism


A = Angle of the Prism


Here, We are working on two prisms.


"d_1 =" Deviation produced by the first Prism


"d_2 =" Deviation produced by the second Prism


"A_1 =" Angle of the first Prism = "5^{o}"


"A_2 =" Angle of the second Prism


"\mu_1 ="  Refractive index of the first Prism = 1.56


"\mu_2 =" Refractive index of the second  Prism = 1.7


Without deviation, means,

"d_1 = d_2 = 0"

Th equations

"d_1 = (\mu_1 - 1) A_1"


"d_2 = (\mu_2 - 1) A_2"


"(\mu_1 - 1) A_1 =(\mu_2 - 1) A_2"

PLug the values in this,

"(1.56 - 1) 5 = (1.7 - 1) A_2"



"A_2 = \frac {0.56 \times 5} {0.7} = 4^{o}"

Answer: Angle of the second Prism

"A_2 = 4^{o}"